OpenStudy (anonymous):

Johnna’s reading rate is 150 words per minute. She is reading a 364-page book that has about 275 words on each page. At this rate, which is closest to the amount of time Johnna should take to read the book? A 10 hrs B 11 hrs C 12 hrs D 13 hrs

OpenStudy (texaschic101):

find out how many words are in the book by multiplying 364 x 275. Then divide that by 150 and that will tell you how many minutes it takes to read the book. Then divide that by 60 and that will tell you how many hours it takes to read the book.

OpenStudy (anonymous):

We first need to find the total number of words in the book. We can do this by multiplying the average number of words/page by the total number of pages: \[Words _{total}=\frac{Words}{page} * Number of Pages\] \[Words _{total}=275\frac{Words}{page} * 364 Pages\] \[Words _{total}=100,100 Words\] Now we can figure out how long it takes her to read those words. We do this be dividing the total number of words, by how many words she can read in one minute: \[Time _{total}=\frac{Total Words}{Words/minute}\] \[Time _{total}=\frac{100100Words}{150Words/minute}\] \[Time _{total}=667.33minutes The answer we want is in hours, but we have an answer in minutes, so we need to convert: 667.33mins / (60 mins / hour) = 11.122 hours The closest answer is (B): 11 Hours.

OpenStudy (texaschic101):

I agree....11 hrs

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

thats what we are here for :)